A body of mass 12kg is projected vertically upwards and made a trajectory with base length 8m for 3secs when the body is projected vertically downwards it dropped into a cylinder containing a liquid of relative density 0.92 if the distance traveled by the body vertically downwards is given as y=3t
4
−4t
3
−8
Find
1. The weight of the body
2. The density of the liquid
3.The maximum height of the projected body
4.The angle of projection
1
Expert's answer
2020-01-29T14:12:50-0500

1)Given the mass of the body is 12 kg

then weight ="mg=12*10=120 N"


2)Given relative density of liquid is 0.92

then density of the liquid =0.92*1000 = 920 kg/m3


3) equation of trajectory "y=3t^4-4t^3-8"

the given equation is not correct so we cannot calculate the maximum height from the above equation